OpenStudy (anonymous):

(2, -3) is a solution of the equation 3x - y = 3.

sam (.sam.):

You can verify this by substituting into the equation, 3x - y = 3 3(2) - (-3) = 3 6+3=3 \[9\neq3\] So its not the solution to the equation